# Bidiagonal matrix factorisations related to multiple orthogonal polynomials

### v1

@@ 4.1. Bidiagonal factorisations and connection with branched continued fractions
Figure 1. All $2$-Dyck paths from $(0,0)$ to $(6,0)$ $\implies S_{2}^{\,(2)}\left(\bm{\alpha}\right)=\alpha_{0}^{2}+\alpha_{0}\alpha_{1}+\alpha_{0}\alpha_{2}$.
A partial $r$-Dyck path starting at $(0,0)$ only passes by points $(x,y)\in\mathbb{N}^{2}$ such that $y=x-(r+1)m$, where $m\in\mathbb{N}$ is the number of $r$-falls of the path and $x$ is the total number of steps. We are interested in the following generalisations of the $r$-Stieltjes-Rogers polynomials: the generalised $r$-Stieltjes-Rogers polynomials, denoted by $S_{n,k}^{\,(r)}\left(\bm{\alpha}\right)$ with $n,k\in\mathbb{N}$, count partial $r$-Dyck paths from $(0,0)$ to $\big{(}(r+1)n,(r+1)k\big{)}$, and the modified $r$-Stieltjes-Rogers polynomials of type $j$, denoted by $S_{n}^{\,(r;\,j)}\left(\bm{\alpha}\right)$ with $n,j\in\mathbb{N}$, count partial $r$-Dyck paths from $(0,0)$ to $\big{(}(r+1)n+j,j\big{)}$.
@@
More generally, for any $k\in\mathbb{N}$, let $f_{k}(t)$ be the generating function for partial $r$-Dyck paths which start and end at height $k$ without ever going below it (i.e., $f_{k}(t)$ is $f_{0}(t)$ with each $\alpha_{i}$ replaced by $\alpha_{i+k}$). The sequence $\left(f_{k}(t)\right)_{k\in\mathbb{N}}$ satisfies the functional equation (see [20, Eqs. 2.26-2.27])
@@
Successively iterating the former we find that (see [20, Eqs. 2.28])
@@
The right-hand side of (4.8) is known as a (Stieltjes-type) $r$-branched continued fraction, because it reduces to a classical Stieltjes continued fraction when $r=1$.
@@
Let $\mathrm{H}=\left(h_{i,j}\right)_{i,j\in\mathbb{N}}$ be an infinite matrix with entries in a commutative ring $r$. If all the powers of $\mathrm{H}$ are well-defined (which is the case when $\mathrm{H}$ is an Hessenberg matrix), we can define an infinite matrix $A=\left(a_{n,k}\right)_{n,k\in\mathbb{N}}$ such that $A_{n,k}=\left(\mathrm{H}^{n}\right)_{0,k}$. We call $\mathrm{H}$ the production matrix and $A$ the output matrix.
To prove Theorem 4.1, it is instrumental to recall the following method to construct Stieltjes-type branched continued fractions introduced in [20].

### v2

@@ 2.3. Lattice paths, branched continued fractions, and production matrices
Figure 1. All 22-Dyck paths from $(0,0)$ to $(6,0)$ $\implies S_{2}^{\,(2)}\left(\bm{\alpha}\right)=\alpha_{0}^{2}+\alpha_{0}\alpha_{1}+\alpha_{0}\alpha_{2}$.
A partial $r$-Dyck path starting at $(0,0)$ only passes by points $(x,y)\in\mathbb{N}^{2}$ such that y=x(r+1)my=x-(r+1)m, where $m\in\mathbb{N}$ is the number of $r$-falls of the path and $x$ is the total number of steps. We are interested in the following generalisations of the $r$-Stieltjes-Rogers polynomials (introduced in [22] and [11], respectively):
  • the modified $r$-Stieltjes-Rogers polynomials of type $j$, denoted by $S_{n}^{\,(r;\,j)}\left(\bm{\alpha}\right)$ with $n,j\in\mathbb{N}$, are the generating polynomials of the partial $r$-Dyck paths from $(0,0)$ to $\big((r+1)n+j,j\big)$,
  • the generalised $r$-Stieltjes-Rogers polynomials of type $j$, denoted by $S_{n,k}^{\,(r;\,j)}\left(\bm{\alpha}\right)$ with $n,k,j\in\mathbb{N}$, are the generating polynomials of the partial $r$-Dyck paths from $(0,0)$ to $\big((r+1)n+j,(r+1)k+j\big)$.
@@
More generally, for any $k\in\mathbb{N}$, let $f_{k}(t)$ be the generating function for partial $r$-Dyck paths which start and end at height $k$ without ever going below it (i.e., $f_{k}(t)$ is $f_{0}(t)$ with each $\alpha_{i}$ replaced by $\alpha_{i+k}$). The sequence $\left(f_{k}(t)\right)_{k\in\mathbb{N}}$ satisfies the functional equation (see [22, Eqs. 2.26-2.27])
@@
Successively iterating the former we find that (see [22, Eqs. 2.28])
@@
The right-hand side of (2.23) is known as a (Stieltjes-type) $r$-branched continued fraction, because it reduces to a classical Stieltjes continued fraction (also known as an S-fraction) when $r=1$.
@@
Let $\mathrm{H}=\left(h_{i,j}\right)_{i,j\in\mathbb{N}}$ be an infinite matrix with entries in a commutative ring $R$. If all the powers of $\mathrm{H}$ are well-defined (which is the case when $\mathrm{H}$ is an Hessenberg matrix), we can define an infinite matrix $\mathrm{S}=\left(s_{n,k}\right)_{n,k\in\mathbb{N}}$ by $s_{n,k}=\left(\mathrm{H}^{n}\right)_{0,k}$ for all $n,k\in\mathbb{N}$. We call $\mathrm{H}$ the production matrix and $\mathrm{S}$ the output matrix.
3. Bidiagonal factorisations linked to the Gauss-Borel factorisation
For a system of linear functionals $\left(v_{1},\cdots,v_{r}\right)$, we consider its Christoffel transformations
$V^{[j]}=\left(v_{j+1},\cdots,v_{r},x\,v_{1},\cdots,x\,v_{j}\right),\;0\leq j\leq r.$ (3.1)
In particular, $V^{[0]}=\left(v_{1},\cdots,v_{r}\right)$ and $V^{[r]}=\left(x\,v_{1},\cdots,x\,v_{r}\right)$.
In this section, we show that the existence and uniqueness of the multiple orthogonal polynomials on the step-line with respect to all systems $V^{[j]}$ in (3.1) is a sufficient condition to obtain bidiagonal factorisations of the Hessenberg recurrence matrices of these systems, with all nontrivial entries of the bidiagonal matrices different from zero. In particular, we obtain a factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ of the Hessenberg recurrence matrix of the system $\left(v_{1},\cdots,v_{r}\right)$.
We start by using the Gauss-Borel factorisation of the moment matrices of the systems $V^{[j]}$ to express the nontrivial entries of the bidiagonal matrices appearing in our factorisations. Then, we rewrite the formulas for the nontrivial entries of the bidiagonal matrices, firstly using coefficients of the type I and type II multiple orthogonal polynomials on the step-line with respect to the systems $V^{[j]}$ instead of their Gauss-Borel factorisation, and later using the determinants of the finite leading principal submatrices of the moment matrices of the systems $V^{[j]}$.
The following theorem is the main theoretical result of this section, where we show that the existence and uniqueness of the multiple orthogonal polynomials on the step-line with respect to all systems $V^{[j]}$ in (3.1) is a sufficient condition to obtain bidiagonal factorisations of the Hessenberg recurrence matrices of these systems, using the entries of either one of the triangular matrices appearing in the Gauss-Borel factorisation of the moment matrices of the systems $V^{[j]}$.
Theorem 3.1.
For a system of linear functionals $\left(v_{1},\cdots,v_{r}\right)$ such that the multiple orthogonal polynomials on the step-line with respect to all systems $V^{[j]}$, $0\leq j\leq r$, in (3.1) exist and are unique, let ${\mathrm{M}^{[j]}=\mathrm{C}^{[j]}\mathrm{D}^{[j]}}$ be the Gauss-Borel factorisation of the moment matrix $\mathrm{M}^{[j]}$ of the system $V^{[j]}$ for each $0\leq j\leq r$, where $\mathrm{C}^{[j]}=\left(c^{[j]}_{i,k}\right)_{i,k\in\mathbb{N}}$} and $\mathrm{D}^{[j]}=\left(d^{[j]}_{i,k}\right)_{i,k\in\mathbb{N}}$} are, respectively, a unit-lower-triangular matrix and a non-singular upper-triangular matrix. Then, for any $0\leq j\leq r$, the Hessenberg recurrence matrix $\mathrm{H}^{[j]}$ of the system $V^{[j]}$ admits the bidiagonal factorisation
where $\mathrm{L}_{1},\cdots,\mathrm{L}_{r}$, and $\mathrm{U}$ are bidiagonal matrices of the form in (1.2), with nontrivial entries $\left(\alpha_{m}\right)_{m\in\mathbb{N}}$ all different from zero such that, for all $n\in\mathbb{N}$,
$\alpha_{(r+1)n}=\frac{d^{[r]}_{n,n}}{d^{[0]}_{n,n}}=c^{[0]}_{n+1,n}-c^{[r]}_{n,n-1},\quad\text{with }c^{[r]}_{0,-1}=0,$ (3.3)
and
$\alpha_{(r+1)n+i}=\frac{d^{[i-1]}_{n+1,n+1}}{d^{[i]}_{n,n}}=c^{[i]}_{n+1,n}-c^{[i-1]}_{n+1,n}\quad\text{for }1\leq i\leq r.$ (3.4)
To prove this theorem, we need the following auxiliary result.